Run Coordination Meetings: Spring run

From GlueXWiki
Jump to: navigation, search

General information

Weekly Meetings, 8:45am, Wednesdays, Hall D conference room. Held by the Run Coordinator. Present Run Coordinator: Alexandre Deur

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1]

2.) To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

4.) More information on connecting to bluejeans is available. The old ESNET and SeeVough Information is archived here.

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Meeting General Agenda

March 12th 2015 meeting

  1. Report relevant updates from 7:50 and 8:00 daily MCC meetings:
    • Accelerator expects to start delivering beam to Hall D next Monday, March 9th.
    1. Apr. 3 - May 4: Physics run (Halls A/B/D).
  2. Report on Work in the Hall:
    • If you have work to do in the Hall, please notify Tom (Hall work coordinator) and if necessary, document it in the electronic logbook.
    • OSL need to be worn everywhere in Hall D.
  3. Hot Check OUT Status 100% ready ; 0% checked ; 0% not ready
  4. Spring 2015 runplan

Installation punch list

(items in red: must be done by 02/27/15 ; Blue: must be done by 03/03/15 ; Green: Nice to have it

Reset HCO

LH2 Target

  1. Installation Tim. Now on the platform. Survey and alignment were finalized today. New flange for the ST done. ST is installed. The target is 18mm upstream compared to ST. ST is in the correct position. We will run that way for spring.
  2. </span> Controls Target Group. Need to assign a dedicated screen for target control. On track.
  3. Install a 1 mm solid CH2 target on the nose of the start counter


  1. Build and install a frame at JLab for testing controls/tune procedure - Tim.
  2. Install frame with test diamonds and foils. There will be three test diamonds (J1a50, J2a100 S145-S90), two 10 μm Al foils, and one 100 μm Al foil with a holed 10 μm Al foil pressed on it (3mm hole). .
  3. Richard Jones will provide the diamonds for initial tests. They will be shipped in 2 weeks- see GlueX diamond inventory
  4. Controls (Hovanes)
  5. Include Status and Controls on the general EPICS GUI (Hovanes)
  6. Port and test the tuning software: Ken L. will come on Apr. 17th but will work remotely on importing his program. Paul M. is his JLab contact.Paul will take care of the initial tests, with Franz K. and Eugene Pasyuk support.
  7. Install nominal diamonds (sometimes in April (1 shift of work)).

Active collimator (remains in its current position in front of 5mm hole and attached to primary collimator)

  1. Test installation in front of 3.4 mm hole. (Time needed for the process:2×15min.)
  2. Check alignment with pin for 5mm and 3.4mm holes
  3. Check noise with new cable configuration
  4. Alex B. found some noise that was already present in Fall. The present situation is thus similar to the Fall run.

Upstream profiler installed on a remotely controlled insertable holder

  1. <strike>Design/construction: Tim. Parts are ordered. They will arrive at the end of Feb. ⇒ Profiler may initially be static.
  2. Installation/Control . Tom/Hovanes. 1 to 2 weeks to install and test controls once the parts arrive.
  3. Survey. Done (Static config.)

Survey Tim. 6h

  1. Realign photon beam line after bellow installation
  2. P.S. detectors
    • coarse detectors
    • high granularity detectors will not be surveyed
  3. Total Absorption Counter (TAC)
  4. LH2/start counter
  5. Diamonds
  6. Photon Polarimeter (will not be surveyed this time)
  7. Solenoid (in case of floor settling)
  8. Upstream profiler (Static config.)
  9. CDC

Solenoid George/Beni.

  1. Power supply maintenance: Danfysik Technician starts Jan 22, finish by Feb 1
  2. Problematic connectors: done
  3. <strike>Power Supply Robustness</strike>. Identify available spares
  4. New VCL Flow Controllers: Vendor to supply new controllers Should be brought to JLab on Friday. Will be installed Monday (~2h)
  5. Cryo trips: Fixed; chart recorder reading flat-lines at 24V
  6. Awkward ground fault fuse check: Requirement eliminated, simplify using high amp fuse by Feb 6
  7. Annual system maintenance: Complete, all power connections and control terminations re-torqued.
  8. Simplified Ramp- up/ down documentation.
  9. Defective He Level control (Cryo Group issue).

Start Counter Eric. P.

  1. Remove the piece of tape in the beam path (to be done while the solid target is installed, ). (Eric P.)
  2. Check channels 1,2,22 and 23. </strike> Make sure cables are not swapped (Eric P.)
  3. Install 1mm CH2 solid target on the nose
  4. Reinstallation
  5. Recabling
  6. ST rotation: ST appears to be rotated by 11.25°. We will run that way.

FDC Lubomir

  1. Chiller is back. Need to be checked re-installed.
  2. Implement power-off switch for chiller safety
  3. (Also for CDC) Analysis of Ar/CO2 gas mixture. Sensor Installation.
  4. Order spare chiller (Tim)

CDC Beni.

  1. All channels are now working (Beni)
  2. Cable re-routing Fernando.
    • Rerouting of LV cables is done.
    • Rerouting of HV cables is done. Will be tested next week.
  3. Ready for cosmic run (BCal trigger) Replaced a few ADC already. Codi put new firmware in ADC.
  4. The problem of ADC crashing when 2 events come in too close in time is fixed
  5. Installation of interlock in case air cooling stops. Dave B.
  6. The major gas leakage that started on Monday 2/9/15, 6pm is fixed
  7. Adjust the gas mixture with less argon.

Pair Spec. Alex S.

  1. Check 8 channels Fernando/Alex S./Vlad. Arm A: 2 channels to be looked at. Arm B: 6 channels. Will check cables but will not do anything beyond (too risky, given the time left before the run)
  2. User-friendly GUI allowing all time monitoring of coinc. rates is ready Alex S.

FCal Hovanes

  1. Voltage GUI
    • Improve underlying control software. Hovanes is working on it.
    • Improve the GUI. Manuel.
    • Status of the noisy channels ? Will be dealt with trigger scheme (see Trigger section).

BCal Elton

  1. Improvement of the air dryness (N2 flow increases)
  2. DAQ check with BCal Turned on pulser system but could not do the check due to networking problems.


  1. Voltage Range Fernando. Done: changed from 1V to 2V.
  2. Problem with discriminator crate.


  1. Tagger
    • Improve shielding at entrance of tagger magnet. Done
    • shield e- beam flange. Done
      • Shield the bottom of the photon beam line pipe. Done
      • Use "Shadow shield (raise current wall just downstream of tagger magnet). Done
      • Add to the green wall a layer of 4 of polyethylene in the sheet of flame ( from beam line to south wall). Done
      • Fill empty beam-pipe hole in the green wall with polyethylene beads. Done.
    • Shield near tagger rack
  2. Collimator cave: shield gamma probe Done

Radiation monitors

  1. In order
  2. Install new radiation monitors
  3. Update readout Hovanes, Pavel D., Dan Sexton
  4. Lay cables for 4 new radiation monitoring probes. (Fernando)

Trigger Alex S.

  1. Reduction of switching noise in FCal trigger Alex S.
    • Implement a trigger with high threshold for high energy signals and with SC (or TOF) coincidence for low energy signals.
  2. Trigger connection to TDC (Beni) Cables made by Chris for redundant timing information. Almost done. Waiting on hardware. Partially tested. Seems good so far.
  3. Firmware update: Put latest version.
  4. Change jumper in distribution board. Done. Will run BCal to check that everything works fine.</span>
  5. Define trigger types and strategy for next run. Will require coincidence with FCal.
  6. Improve software for scaler readout (Paul M.) Changing method to send data to epics: Run epics IOC on each crates. Should be faster. Will help Hovanes to write necessary software. Should be ready in ~2 weeks.

DAQ: Multiblock mode run at ≈20 kHz Sergey

  1. Problem with FADC. Cody is working on a workaround to be able to read at 5 kHz at ? deadtime
  2. Fix F125 firmware problem Cody is working on fixing it (for next year run)
  3. Run Condition database RCDB interface with DAQ. Done
  4. Need to check new CODA version from CODA group. Need to test it.
  5. DAQ not yet tested on Full scale.
  6. Working on pedestals for ADC

Online David L.

  1. Fix high rate freeze of monitoring histograms Working on it. The problem is identified and a fix is being worked out.
  2. Upgrade O.S. (Red Hat 7) Deferred to after Spring run, except for GLUON01 that is upgraded. Doesn't work when trying to run CSS).
  3. Pedestal stability monitoring/temperature stability (Marc D.)

Accelerator RF timing (Elton/Fernando)

  1. Electronic to convert fiber signal to electronic signal. Done
  2. Fibers from accelerator are installed (Hall + Tagger). Done
  3. Working on the two signal distribution boards. Done
  4. Ordering cables for signal distribution. Done
  5. Everything should be completed by Feb 15th.
  6. Check that boards are not making crates to hang
  7. See spurious pulse outputs from distribution boards, This happens only when the boards are unplugged. Boards need to be redesigned for later runs.

Photon polarimeter (Kei)

  1. Should arrive between mid-February and early March
  2. Need to install a test setup outside the Hall
  3. Fiducialize (outside Hall)
  4. Installation: 1 day
  5. Testing
  6. Survey
  7. Control


  1. Survey is done
  2. Signals are checked
  3. Need to figure out TAC trigger. Alex S. will set-up a self trigger.

Phone Sprint repeater and WiFi.

  1. Will be tested mid-March, when all equipments/controls are ready
  2. Need to let IT know of the test, for safety purpose (won't be able to call 911/firefigheters/guards from cell phones in the hall) . IT have administration authority to turn repeater on/off. IT contact: Brian Hess.
  3. Tom to be the interface with IT.

== Specific safety notice (from Tom/Tim) ==